OpenStudy (anonymous):

Which expression shows the value of sec x°? adjacent ÷ opposite hypotenuse ÷ opposite hypotenuse ÷ adjacent adjacent ÷ hypotenuse

OpenStudy (anonymous):

adjacent/hypo

OpenStudy (anonymous):

sryyyyyyyy

OpenStudy (anonymous):

hypo/adj

OpenStudy (anonymous):

pelletttttt

OpenStudy (anonymous):

shhhiiiitttt

OpenStudy (anonymous):

\(\huge sec x =\frac{1}{cos x}=\frac{1}{\frac{adjacent}{hypo}}=\frac{hypo}{adjacent} \)

OpenStudy (anonymous):

yapppp

OpenStudy (anonymous):

inverse of cosx

OpenStudy (anonymous):

sryyyyy slip of hand